the perimeter of a rectangular room is 100 feet.find the length and width of the room if the length is 2 feet more than 3 times the width.
Answer by checkley77(12844)

L=3W+2
2L+2W=100
2(3W+2)+2W=100
6W+4+2W=100
8W=100-4
8W=96
W=96/8
W=12 FT. ANS. FOR THE WIDTH.
L=3*12+2=36+2=38 FT. ANS. FOR THE LENGTH.
PROOF:
2*38+2*12=100
76+24=100
100=100